OpenStudy (anonymous):

Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true. Expression #1 Expression #2 (y9)(2y2)3 (9y)(2y3)2 The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (cwrw238):

using the laws of logarithms y^9* (2y^2)^3 = y^9 * 2^3* y^(2*3) = y^9 * 8 * y^6 = 8 y^15 - can you simplify expression 2 in same way?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

2 simplifies into 1

OpenStudy (cwrw238):

right good job - so 3rd option is correct

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw